# Double strangeness pentaquarks and other exotic hadrons in the $\Xi_{b}\to\Xi\ J/\Psi\ \phi$ decay

### v1

@@ V.1 The $J/\psi~{}\phi$ mass distribution
We start studying the $J/\psi~{}\phi$ mass distribution, which will allow us to see the $X(4140)/X(4160)$ interplay. We consider two possibles models: in the first one, which is based in the latest experimental results (Refs. LHCb:2016axx ; LHCb:2016nsl ), we assume existence of only one broad $X(4140)$ resonance with a pole position of $M_{R}=4146.5~{}{\rm MeV}$ and a width of $\Gamma=83~{}{\rm MeV}$; while the second model assumes the existence of a narrow $X(4140)$ plus a wide $X(4160)$ resonance, based on earlier experimental results Wu:2010vk ; Hofmann:2005sw ; Yuan:2012wz ; Xiao:2013yca ; Garcia-Recio:2013gaa and a recent theoretical model Wang:2017mrt .
We start with the model with a wide $X(4140)$. The corresponding results are shown in Fig. 22, where we can see the $J/\psi~{}\phi$ invariant mass distribution for the tree level calculations (Fig. 19) and the corresponding spectrum when we also take into account the FSI (Figs. 21). As we can see, the effect of the FSI is rather small. This will be also true for the model with two $X$ resonances, so in this section we will further present only the results that include the FSI.
@@
Once seen the results for the one resonance model we are going to discuss the results generated with two resonances, where similarly to the results of Ref. Magas:2020zuo we adjust the parameter $\beta$ to a value which generates two $X$ resonances of a similar strength. The $X(4140)$ resonance has a mass of $M_{R}=4132~{}{\rm MeV}$ and a width of $\Gamma=19~{}{\rm MeV}$, while the $X(4160)$ can be generated as $D^{*}_{s}\bar{D}^{*}_{s}$ molecule, as we explained in Section IV.2. The corresponding model contains a loop-function that diverges and must be regularised. In Ref. Molina:2009ct this loop function is regularised using a dimensional regularisation method but, as we proved earlier, such a scheme can produce fake poles, so in this work we use the cut-off regularisation method with $\Lambda=650~{}{\rm MeV}$, which reproduces the results obtained using the dimensional regularisation from Molina:2009ct .
The $J/\psi~{}\phi$ invariant mass distribution for two $X$ resonances can be seen in Fig. 23, where we can see a narrow $X(4140)$ peak at $4132~{}{\rm MeV}$ and a rather wide $X(4160)$ peak. Note that the $X(4160)$ also produces a cusp around $4224~{}{\rm MeV}$, which corresponds to the $D^{*}_{s}\bar{D}^{*}_{s}$ threshold. As it is mentioned in Ref. Magas:2020zuo , if this cusp is experimentally detected it would strongly suggest a molecular interpretation of $X(4160)$, as well as the existence of a narrow $X(4140)$ and a wide $X(4160)$ resonance in this energy region.
@@
For the results presented in Fig. 23 the parameter $\beta$, Eq. 33, is taken as $\beta=\beta_{0}/2.6$, where $\beta_{0}$ is the value taken in Ref. Wang:2017mrt . We would also like to analyse the sensibility of the obtained distribution with respect to the parameter $\beta$ and, for this purpose, we generated similar spectra for different values of $\beta$ and present these in Fig. 24. As it can be expected from eq. (33), changing the values of $\beta$ mainly affects the height of the $X(4140)$ peak.
Figure 24: The solid lines represent the $J/\psi~{}\phi$ mass spectrum generated using the $X(4140)$ and the $X(4160)$ resonances for different values of the weight factor $\beta$, which controls the relative contributions of the two resonances; $\beta_{0}$ is the value of $\beta$ given in Ref. Wang:2017mrt
@@ V.2 The $J/\psi~{}\Xi$ mass distribution
Now we discuss the results for the $J/\psi~{}\Xi$ mass distribution, which is the channel where we expect to see the signature of the $S=-2$ pentaquark. The properties of these pentaquark, dynamically generated by our model, are presented in Tab. 7. This state has a mass $M_{R}=4633.38~{}{\rm MeV}$ and a width $\Gamma_{R}=79.58~{}{\rm MeV}$, and couples rather strongly to the $J/\psi~{}\Xi$ channel, with a coupling constant value of $g_{J/\phi\Xi}=-1.62+0.38i$. We will take these values for the pole position and coupling as the nominal ones and we will vary them within a reasonable range to explore the sensitivity of our results with respect to the pentaquark parameters. We also will study the effect of narrow state in the $J/\psi~{}\Xi$ channel by using the parameters obtained in ref. Roca:2024nsi , where the state has a mass $M_{R}=4617~{}{\rm MeV}$, a width $\Gamma_{R}=12~{}{\rm MeV}$ and the coupling constant to the $J/\psi~{}\Xi$ channel is $g_{J/\phi\Xi}=0.66+0.0i$.
Figure 25: The $J/\psi~{}\Xi$ spectrum, where the purple line corresponds to the tree level diagram and the green and red one is obtained also taking into account the final state interactions that generates the $S=-2$ pentaquark which has $M_{R}=4633~{}MeV$ and $\Gamma=79~{}MeV$ or $M_{R}=4617~{}MeV$ and $\Gamma=12~{}MeV$ respectively.
Proceeding in a similar manner as in the previous section, we start showing our results computed with one wide $X(4140)$ resonance. In Fig. 25 we can see the $J/\psi~{}\Xi$ spectrum, where the solid line is obtained using only the tree-level diagram, while the dashed lines include also the FSI effects. In these latter curves we should see the signal of the $S=-2$ pentaquark, generated in the $J/\psi~{}\Xi$ interaction. First of all, as can be seen in Fig. 25, the background shows itself a peak structure around $4730~{}{\rm MeV}$. The pentaquark, in both studied cases, interferes with the background positively and has an important effect on the spectrum due to its strong coupling to the $J/\psi~{}\Xi$ channel. The presence of the pentaquark in FSI manifest itself as the appearance of a bump around the nominal pentaquark mass of $4633~{}{\rm MeV}$ in the case of the wide resonances from our model. That bump is not very prominent, however it seems to be detectable. Meanwhile for the case we have a narrow pentaquark from Roca:2024nsi the peak that is formed due to its existence is not much higher than the bump and also perfectly detectable. The explanation for this phenomenon is rather straightforward: narrow pentaquark is much sharper and at the first glance seems to be easier detectable, however it is narrower, because its coupling to the lowest open channel $J/\psi\Xi$ is smaller, and thus the amplitude $T_{J/\psi\Xi,J/\psi\Xi}\sim g_{J/\psi\Xi}^{2}$ is also smaller, producing at the end of the day approximately the same signal to noise ration.
@@
Figure 27: The purple line dashed represents the $J/\psi~{}\Xi$ spectrum computed with a narrow $X(4140)$ plus a $X(4160)$ having a resonance with $M_{R}=4633~{}MeV$ and $\Gamma=79~{}MeV$, while the green and the blue ones show similar results but having a resonance with, respectively $M_{R}=4617~{}MeV$ and $\Gamma=12~{}MeV$.
Let’s now consider the calculation with a narrow $X(4140)$ and a wide $X(4160)$ resonances. The results (with FSI) can be seen in Fig. 26, where the green and the blue curves show the individual contributions of the $X(4140)$ and the $X(4160)$ resonances respectively, while the purple solid curve is the sum of both contributions. Note that the $X(4160)$ contribution is dominant with respect to that of the $X(4140)$. It is also important to see that, in contrast to the previous case, the signal of the pentaquark interferes negatively with the background. This can be seen more clearly in Fig. 27, where we compare the contributions from wide and narrow pentaquarks. Again we see that the presence of $P_{css}$ is well detectable in ambos cases, and that both produce approximately the same signal to noise ration. For the further studies we will concentrate on wider pentaquark produced by our model, as discussed in section III.2.
In Fig. 28 where we compare the contribution from one and two $X$ resonance models with FSI, varying the mass of the wide pentaquark and its coupling to the $J/\psi~{}\Xi$ channel. We do that since we know there can exist some theoretical uncertainties, so we modify the pentaquark pole position by ±$\pm 50~{}{\rm MeV}$ and we allow for a $20\%$ of variation of the coupling of the pentaquark to the channel $J/\psi~{}\Xi$. The results presented in Fig. 28 clearly show that the presence of the double strangeness pentaquark is experimentally detectable in all cases. Actually, in the case of one wide $X(4140)$ this task seems to be much simpler. And certainly the experimental measurement of the $J/\psi~{}\Xi$ invariant mass spectrum from the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay would also allow one to distinguish between one wide $X(4140)$ or two $X$ models.
@@ V.3 The $\phi~{}\Xi$ mass distribution
Finally, we present the results for the $\phi~{}\Xi$ invariant mass distribution where we expect to see the signal of the $\Xi(2500)$. This state appears in the PDG PDG as a one-star state, with a mass of $M_{\Xi(2500)}=2500~{}{\rm MeV}$ and there exist two experimental groups that reported different values for its width, namely $150~{}{\rm MeV}$ Alitti:1969rb and $50~{}{\rm MeV}$ Aachen-Berlin-CERN-London-Vienna:1969bau . Based on this information we decided to generate the spectrum for tree different cases, $\Gamma=50,~{}100,~{}150~{}{\rm MeV}$. We do not know either which is the coupling constant of this state to the $\phi~{}\Xi$ channel, therefore we explore a range of values for this coupling, which would allow us to see the signal of the $\Xi(2500)$ in generated the spectrum.
@@
The corresponding results are the presented in Fig. 29. The value of the coupling constant, needed to generate a significant interference with the background, as we could expect, is larger if the state has a larger width. We can see that if $\Gamma=50~{}{\rm MeV}$ then the minimum absolute value of the coupling is around $0.7$, while for a $\Xi(2500)$ with $\Gamma=150~{}{\rm MeV}$ this value is around $1.7-1.9$. Similarly to the previous case, the interference of the resonance with the background is positive for the case of one wide $X(4140)$ resonance, while for the case of a narrow $X(4140)$ and wide $X(4160)$ resonances the interference is negative. Thus, we see that studying this spectrum would also be very interesting if the coupling of the $\Xi(2500)$ to the $\phi~{}\Xi$ states is large enough.
@@ VI Conclusions
The discovery in the last years of many exotic hadrons not fitting into the conventional quark composition model, has converted this topic into an important field of research. Motivated by this fact, in the present work we studied some resonances that could be interpreted as quasi-bound states of an interacting meson-baryon pair in the strangeness $-2$ and isospin $1/2$ sector, employing effective Lagrangians for describing the exchange of a vector meson in the t-channel. We also proposed a decay process $(\Xi_{b}\to\Xi~{}J/\psi~{}\phi)$ suitable to detect some of those exotic hadrons in all two-particle channels.
@@
First of all, in this work we find two poles, which can be associated to the $\Xi(1620)$ and $\Xi(1690)$ resonances from PDG PDG , in the scattering amplitude of the interaction of pseudoscalar mesons with baryons in the light-channel sector. By modifying some parameters in a reasonable range we can reproduce the experimental masses of the $\Xi(1620)$ and $\Xi(1690)$, producing two states at $1600~{}{\rm MeV}$ and $1685.40~{}{\rm MeV}$. With this result we can conclude that the $\Xi(1620)$ and the $\Xi(1690)$ may have a meson-baryon molecular origin. The $\Xi(1620)$ would have an important $\pi\Xi$ component, around $50\%$, with a $20\%$ mixture of $\bar{K}\Lambda$. The $\Xi(1690)$ would be mainly a $\bar{K}\Sigma$ molecule with a small component $(2\%)$ of $\eta\Xi$. Both of these resonances have well defined spin-parity $(J^{\pi}=\frac{1}{2}^{-})$, and thus a measurement of the spin-parity of these states could confirm or refute this hypothesis.
@@
For the light-channel sector of the VB interaction we can also find two resonances in the scattering amplitude, which could be associated with two states listed in the PDG, namely the $\Xi(1820)$ and $\Xi(1950)$ PDG . Similarly to what we did in the PB section, by modifying some parameters we can bring the positions of these two states closer to their experimental mass, but the theoretical widths are much smaller than the experimental ones. This might be solved by allowing the coupling between the vector mesons with the pseudoscalar mesons, and thus opening decays to lighter channels and correspondingly increasing the width of the dynamically generated states. If the present interpretation is assumed to represent the nature of these states, we could understand the $\Xi(1820)$ as a mixture of $\phi\Xi$ $(33\%)$ and $\bar{K}^{*}\Lambda$ $(13\%)$ components, while the $\Xi(1950)$ would be a mixture of $\bar{K}^{*}\Sigma$ $(50\%)$, $\omega\Xi$ $(10\%)$ and $\phi\Xi$ $(10\%)$ ones. Note that these states are degenerate in spin-parity and form a $J^{\pi}=\frac{1}{2}^{-}$, $J^{\pi}=\frac{3}{2}^{-}$ doublet. This would be compatible with the observed quantum numbers of the $\Xi(1820)$, which are $J^{\pi}=\frac{3}{2}^{-}$.
@@
Alternatively our model allows to generate these two resonances in the region of the $\Xi(1950)$, producing an apparent wide resonance. This is a possibility mentioned in the PDG, where it is said that the $\Xi(1950)$ may in fact be more than one $\Xi$ state.
@@
Furthermore, the results for the heavy channels are quite encouraging, as they point towards the existence of doubly strangeness pentaquarks with hidden charm, which can be generated thanks to the coupled-channel structure of our model. The absence in this sector of a long range interaction mediated by pion-exchange also makes the search for these states specially interesting. If they do exist, their interpretation as molecules would require a change of paradigm, since they could only be bound through heavier-meson exchange models as the one employed in this work, see ref. Marse-Valera:2022khy for more details.
@@
Indeed, in the pseudoscalar meson baryon interaction sector we can generate one hidden charm $S=-2$ baryon with $M=4493~{}{\rm MeV}$ and $\Gamma=73~{}{\rm MeV}$, which can be interpret as a $\bar{D}\Omega_{c}$ molecule, and the quantum numbers of this state would be $J^{\pi}=\frac{1}{2}^{-}$. In the vector meson baryon interaction sector we can also generate one state with $M=4633~{}{\rm MeV}$ and $\Gamma=79~{}{\rm MeV}$ and, similarly to the previous case, this state can be understood as a $\bar{D}^{*}\Omega_{c}$ molecule which would be degenerate in spin-parity forming a $J^{\pi}=\frac{1}{2}^{-}$, $J^{\pi}=\frac{3}{2}^{-}$ doublet.
@@
The similar model of Ref. Roca:2024nsi also find such states, practically at the same mass, but with much narrower width. In the further study, Ref. Oset:2024fbk , the authors suggest the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decays in order to observe the $P_{css}$ with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$ generated via pseudoscalar-baryon interaction. In contrast, in the present manuscript we have concentrated at the possibility to detect potential $P_{css}$ state generated in VB interaction in $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay.
@@
Another motivation to study the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay came form the Dalitz plots of Fig. 15, which proved the potential possibility to study several exotic states that may be detected in all three two-body invariant mass spectra. Similarly to what is done in Magas:2020zuo , we develop two different models for the interpretation of the $X(4140)$ and $X(4160)$ resonances. The first model considers that there only exists one wide $X(4140)$ state while the other claims the existence of a narrow $X(4140)$ resonance plus a $X(4160)$ one. Consequently in the $J/\psi~{}\phi$ mass spectrum, due to the nature of the $X(4160)$ as a $D^{*}_{s}\bar{D}_{s}^{*}$ Molina:2009ct bound system, there appears a cusp near the threshold of this channel $(4224~{}{\rm MeV})$. Therefore, if this cusp is detected, it would be a clear signal that the two-resonances model is the one that represents the nature of the $X$ resonances, and also that the $X(4160)$ is a molecular state.
@@
The $X(4274)$, $X(4350)$ and the $X(4500)$ resonances may also leave a signal at higher $J/\psi$ invariant masses in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay, but in this work these states where not included as we do not have a theoretical model that generates them.
@@
In the $J/\psi~{}\Xi$ spectrum, we studied the possibility of detecting the doubly strange pentaquark with hidden charm, which we introduce as a Breit-Wigner using the parameters obtained from the coupled channel approach developed in this work in section III.2. This permits us to keep the model simple and do variations on the parameters easily. The results are very promising, since if the mass of this state lies in the region $4580-4680~{}{\rm MeV}$ it has a good chance to be experimentally detected. We have also shown that in the case of a single wide $X(4140)$ model the pentaquark signal interferes positively with the background, while for the two $X$ model this interference is negative. This potentially allows one to differentiate between these two situations once the corresponding spectrum will be measured.

### v2

@@ V.1 The $J/\psi~{}\phi$ mass distribution
We start studying the $J/\psi~{}\phi$ mass distribution, which will allow us to investigate the $X(4140)/X(4160)$ interplay. We consider two possibles models: the first one, which is based in the latest experimental results [56, 66], assumes the existence of only one broad $X(4140)$ resonance with a pole position at $M_{R}=4146.5~{}{\rm MeV}$ and a width of $\Gamma=83~{}{\rm MeV}$, while the second model assumes the existence of a narrow $X(4140)$ plus a wide $X(4160)$ resonance, based on earlier experimental results [14, 15, 16, 17, 18] and a recent theoretical model [62].
We start with the model with a wide $X(4140)$. The corresponding results are shown in Fig. 21, where we can see the $J/\psi~{}\phi$ invariant mass distribution for the tree level calculations (see the diagram of Fig. 18) and the corresponding spectrum when we also take into account the final-state interaction (FSI) diagrams of Fig. 20. As we can see, the effect of FSI is rather small. This will be also true for the model with two $X$ resonances, so in this section we will further present only the results that include the FSI effects.
@@
We next discuss the results generated with two resonances, where, similarly to the results of Ref. [61], we adjust the parameter $\beta$ to a value which provides a similar strength for the contribution of the two $X$ resonances. The $X(4140)$ has a mass of $M_{R}=4132~{}{\rm MeV}$ and a width of $\Gamma=19~{}{\rm MeV}$, while the $X(4160)$ can be generated as a $D^{*}_{s}\bar{D}^{*}_{s}$ molecule, as we explained in Section IV.2. The corresponding model contains a loop-function that diverges and must be regularized. While a dimensional regularization scheme was employed in Ref. [60], we proved earlier that such a scheme can produce fake poles. Therefore, in this work we use the cut-off regularization method with $\Lambda=650~{}{\rm MeV}$, which reproduces the results obtained using the dimensional regularization scheme in [60] in the energy region of interest.
The $J/\psi~{}\phi$ invariant mass distribution for two $X$ model is displayed in Fig. 22, where we can see a narrow $X(4140)$ peak at $4132~{}{\rm MeV}$ and a rather wide $X(4160)$ peak. Note that the $X(4160)$ also produces a cusp around $4224~{}{\rm MeV}$, which corresponds to the $D^{*}_{s}\bar{D}^{*}_{s}$ threshold. As it is mentioned in Ref. [61, 62], if this cusp is experimentally detected it would strongly suggest a molecular interpretation of $X(4160)$, as well as the existence of a narrow $X(4140)$ and a wide $X(4160)$ resonance in this energy region.
@@
For the results presented in Fig. 22 the parameter $\beta$ of Eq. (32) is taken as $\beta=\beta_{0}/2.6$, where $\beta_{0}$ is the value employed in Ref. [62]. We would also like to analyze the sensibility of the obtained distribution with respect to the parameter $\beta$ and, for this purpose, we generated similar spectra for different values of $\beta$ and present these in Fig. 23. As it can be expected from Eq. (32), changing the values of $\beta$ mainly affects the height of the $X(4140)$ peak.
Figure 23: The solid lines represent the $J/\psi~{}\phi$ mass spectrum generated using the $X(4140)$ and the $X(4160)$ resonances for different values of the weight factor $\beta$, which controls the relative contributions of the two resonances; $\beta_{0}$ is the value of $\beta$ given in Ref. [62]
@@ V.2 The $J/\psi~{}\Xi$ mass distribution
Now we discuss the results for the $J/\psi~{}\Xi$ mass distribution, which is the channel where we expect to see the signature of the $S=-2$ pentaquark. As seen in Table 7, the pentaquark generated dynamically from the VB interaction has a mass $M_{R}=4633~{}{\rm MeV}$, a width $\Gamma_{R}=80~{}{\rm MeV}$, and couples rather strongly to the $J/\psi~{}\Xi$ channel, with a value of $g_{J/\phi\Xi}=-1.62+0.38i$. We will take these values for the pole position and coupling as the nominal ones and we will vary them within a reasonable range to explore the sensitivity of our results to the pentaquark properties. We will also study the signal in the $J/\psi~{}\Xi$ distribution from a narrower pentaquark by using the parameters of the state obtained in Ref. [22], namely a mass of $M_{R}=4617~{}{\rm MeV}$, a width of $\Gamma_{R}=12~{}{\rm MeV}$ and a coupling constant to the $J/\psi~{}\Xi$ channel of $g_{J/\phi\Xi}=0.66+0.0i$.
Figure 24: The $J/\psi~{}\Xi$ spectrum with a wide $X(4140)$ resonance: the purple line corresponds to the tree level diagram; the green dash-dotted curve is generated by including the final state interaction effects using the full amplitude $T_{J/\phi\Xi,J/\psi\Xi}$ of the VB molecular model; the blue dotted and the red dashed ones are obtained modeling the $S=-2$ pentaquark with a Breit-Wigner shape with $M_{R}=4633$ MeV and $\Gamma=80$ MeV or $M_{R}=4617$ MeV and $\Gamma=12$ MeV, respectively.
Proceeding in a similar manner as in the previous section, we start by showing our results computed with one wide $X(4140)$ resonance. In Fig. 24 we can see the $J/\psi~{}\Xi$ spectrum, where the solid line is obtained using only the tree-level diagram, while the other lines include also the FSI effects. In these latter curves we should see the signal of the $S=-2$ pentaquark. First of all, we note that the background shows itself as a wide structure around $4730~{}{\rm MeV}$. Both the green and blue lines present results for the molecular pentaquark with the caracteristics found by in our model, however the green dashed-dotted line is obtained using the full $T_{J/\psi\Xi,J/\psi\Xi}$ amplitude, while the blue dotted line is obtained by simulating this amplitude by a Breit-Wigner expression, Eq. (25). The red dashed line is also a result of a simulation with the Breit-Wigner formula, however the parameters of the pentaquark state are taken from Ref. [22].
We observe that in our models the pentaquark interferes with the background positively and generates an important effect on the spectrum due to its sizable coupling to the $J/\psi~{}\Xi$ channel. The pentaquark obtained in the present work manifests itself as a bump around its nominal mass of $4633~{}{\rm MeV}$, not very prominent, but seemingly detectable. The difference between green and blue curves comes from the fact that in the wide energy range the full amplitude may differ substantially from its Breit-Wigner approximation, and it gives the idea about the order of uncertainty which such an approximation may generate.
The narrow pentaquark from [22] interferes with the background negatively and leaves a sharper signal in the $J/\psi~{}\Xi$ invariant mass spectrum with approximately the same signal-to-noise ratio than the wider pentaquark.
@@
Figure 26: The $J/\psi~{}\Xi$ spectrum obtained with a narrow $X(4140)$ resonance plus the $X(4160)$: the purple line corresponds to the tree level diagram; the green dash-dotted curve is generated by including the final state interaction effects using the full amplitude $T_{J/\phi\Xi,J/\psi\Xi}$ of the VB molecular model; the blue dotted and the red dashed ones are obtained modeling the $S=-2$ pentaquark with a Breit-Wigner shape with $M_{R}=4633$ MeV and $\Gamma=80$ MeV or $M_{R}=4617$ MeV and $\Gamma=12$ MeV, respectively.
We now consider the calculation with a narrow $X(4140)$ resonance and a wide $X(4160)$ one. The results including FSI effects can be seen in Fig. 25, where the green and the blue curves show the individual contributions of the $X(4140)$ and the $X(4160)$ resonances, respectively, while the purple solid curve is the sum of both. Note that the $X(4160)$ contribution is the dominant one. It is also important to see that, in contrast to the previous case, the signal of the pentaquark interferes negatively with the background. This can be seen more clearly in Fig. 26, where we compare the spectra obtained assuming a wide or a narrow pentaquark. While the narrow option leaves an apparently better distinguishable signal, the presence of the $P_{css}$ is well detectable in both cases, with a similar signal-to-noise ratio. Again the difference between the green and blue curves, actually more prominent that in case of a wide $X(4140)$ resonance, Fig. 24, comes from the difference between employing the full $T_{J/\phi\Xi,J/\psi\Xi}$ amplitude and its Breit-Wigner approximation.
In the following results we will concentrate on the wider pentaquark, i.e. the one produced by our model and discussed in section III.2.
In Fig. 27 we compare the the $J/\psi~{}\Xi$ spectrum for the one (left panels) and the two (right panels) $X$ resonance models including FSI, but varying the mass of the wide pentaquark and its coupling to the $J/\psi~{}\Xi$ channel, in order to implement the effect of some theoretical uncertainties. We modify the pentaquark pole position by ±$\pm 50~{}{\rm MeV}$ (top to bottom panels) and we allow for a $20\%$ of variation of the coupling of the pentaquark to the channel $J/\psi~{}\Xi$ (different colors in each panel). The results of Fig. 27 clearly show that the presence of the double strangeness pentaquark is experimentally detectable in all cases. Actually, in the case of one wide $X(4140)$ (left panels) this task seems to be more easily achievable. And certainly the experimental measurement of the $J/\psi~{}\Xi$ invariant mass spectrum from the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay would also allow one to distinguish, from the different shape of the pentaquark signal, between the one wide $X(4140)$ resonance model or that assuming two $X$ states.
@@ V.3 The $\phi~{}\Xi$ mass distribution
Finally, we present the results for the $\phi~{}\Xi$ invariant mass distribution where we expect to see the signal of the $\Xi(2500)$. This appears as a one-star state in the PDG [7], with a mass of $M_{\Xi(2500)}=2500~{}{\rm MeV}$, and there exist two experimental groups that reported different values for its width, namely $150~{}{\rm MeV}$ [67] and $50~{}{\rm MeV}$ [68]. Based on this information we decided to calculate the spectra shown in Fig. 28. They correspond to the different interpretation of the $X$ states: one wide $X(4140$) (left panels) or a narrow X(4140) plus a wide X(4160) (right panels). From top to bottom, we represent the spectrum for tree different widths, $\Gamma=50,~{}100,~{}150~{}{\rm MeV}$. Finally, as we do not know either which is the coupling constant of this state to the $\phi~{}\Xi$ channel, we explore a range of values for this coupling (different colors in each panel) to find the minimum value which would allow us to see the signal of the $\Xi(2500)$ in the generated spectrum.
@@
We observe, as expected, that the minimum value of the coupling constant needed to generate a significant interference with the background is larger if the state has a larger width. More specifically, if $\Gamma=50~{}{\rm MeV}$ then the minimum absolute value of the coupling is around $0.7$, while for a $\Xi(2500)$ with $\Gamma=150~{}{\rm MeV}$ this value grows to $1.7-1.9$. Similarly to the previous case, the interference of the resonance with the background is positive for the case of one wide $X(4140)$ resonance, while for the case of a narrow $X(4140)$ resonance plus a wide $X(4160)$ one the interference is negative. Thus, studying this spectrum as a potential way of distinguishing between the two models would also be very interesting, provided the coupling of the $\Xi(2500)$ to the $\phi~{}\Xi$ states is large enough for its signal to be seen in the spectrum.
@@ VI Conclusions
The discovery in the last twenty years of many exotic hadrons not fitting into the conventional quark composition model, has converted this topic into an important field of research. Motivated by this fact, in the present work we studied some resonances that could be interpreted as quasi-bound states of an interacting meson-baryon pair in the strangeness $-2$ and isospin $1/2$ sector, employing effective Lagrangians for describing the exchange of a vector meson in the t-channel. We also proposed a decay process $(\Xi_{b}\to\Xi~{}J/\psi~{}\phi)$ suitable to detect some of those exotic hadrons in all two-particle invariant mass spectra.
@@
First of all, in the light flavor sector, the scattering amplitude for the interaction of pseudoscalar mesons with baryons presents two poles, which can be associated to the $\Xi(1620)$ and $\Xi(1690)$ resonances from the PDG [7] compilation. By modifying some parameters in a reasonable range we can reproduce the experimental masses of the $\Xi(1620)$ and $\Xi(1690)$, producing two states at $1600~{}{\rm MeV}$ and $1685.40~{}{\rm MeV}$. We can therefore conclude that the $\Xi(1620)$ and the $\Xi(1690)$ may have a meson-baryon molecular origin. In our model, the $\Xi(1620)$ would have an important $\pi\Xi$ component, around $50\%$, with a $20\%$ mixture of $\bar{K}\Lambda$. The $\Xi(1690)$ would be mainly a $\bar{K}\Sigma$ molecule with a small component $(2\%)$ of $\eta\Xi$. Note, however, that this composition differs from that of the states found by recent theoretical models that include the NLO terms of the Lagrangian, the widths of which are also more in accordance with the experimental values. This interesting topic deserves further investigations.
@@
For the light-channel sector of the interaction of vector mesons with baryons we also find two poles in the scattering amplitude, which could be associated with two resonances listed in the PDG, namely the $\Xi(1820)$ and $\Xi(1950)$ [7]. By modifying some parameters we can bring the positions of these two states closer to their experimental masses, but the theoretical widths are much smaller than the experimental ones. This might be solved by implementing the coupling between the vector meson channels with the pseudoscalar meson ones, thus opening the decay to lighter channels and correspondingly increasing the width of the dynamically generated states. If the present interpretation is assumed to represent the nature of these states, we could understand the $\Xi(1820)$ as a mixture of $\phi\Xi$ $(33\%)$ and $\bar{K}^{*}\Lambda$ $(13\%)$ components, while the $\Xi(1950)$ would be a mixture of $\bar{K}^{*}\Sigma$ $(50\%)$, $\omega\Xi$ $(10\%)$ and $\phi\Xi$ $(10\%)$ ones. Note that these states are degenerate in spin-parity and form a $J^{\pi}=\frac{1}{2}^{-}$, $J^{\pi}=\frac{3}{2}^{-}$ doublet. This would be compatible with the observed quantum numbers of the $\Xi(1820)$, which are $J^{\pi}=\frac{3}{2}^{-}$.
@@
Alternatively our model allows to generate these two resonances in the region of the $\Xi(1950)$, producing an apparent wide resonance. This is a possibility pointed out in the PDG, where it is mentioned that the $\Xi(1950)$ may in fact be more than one $\Xi$ state.
@@
The results for the heavy sector are quite encouraging, as they point towards the existence of double strangeness pentaquarks with hidden charm, which can be generated thanks to the coupled-channel structure of our model. The absence in this sector of a long range interaction mediated by pion-exchange also makes the search for these states specially interesting. If they do exist, their interpretation as molecules would require a change of paradigm, since they could only be bound through heavier-meson exchange models as the one employed in this work, see Ref. [20] for more details.
@@
Indeed, in the pseudoscalar meson baryon interaction sector we can generate one hidden charm $S=-2$ baryon with $M=4493~{}{\rm MeV}$ and $\Gamma=73~{}{\rm MeV}$, which can be interpret as a $\bar{D}\Omega_{c}$ molecule, and the quantum numbers of this state would be $J^{\pi}=\frac{1}{2}^{-}$. In the vector meson baryon interaction sector we can also generate one state with $M=4633~{}{\rm MeV}$ and $\Gamma=80~{}{\rm MeV}$ and, similarly to the previous case, this state can be understood as a $\bar{D}^{*}\Omega_{c}$ molecule which would be degenerate in spin-parity forming a $J^{\pi}=\frac{1}{2}^{-}$, $J^{\pi}=\frac{3}{2}^{-}$ doublet.
@@
The recent similar model of Ref. [22] also finds such states, practically at the same mass, but with much narrower width, a fact tied to the smaller size of the coupled-channel coefficients. In a subsequent study, Ref. [27], the authors suggest the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decay processes in order to observe the $P_{css}$ with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$ generated via a pseudoscalar-baryon interaction. In contrast, in the present manuscript we have focused on the possibility to potentially detect a $P_{css}$ state generated from a vector-baryon interaction in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay process.
@@
Another motivation to study the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay process emerged from the inspection of the Dalitz plots of Fig. 14, which proved the potential possibility to study several exotic states that may be detected in all three two-body invariant mass spectra. Similarly to what is done in [61], we have developed two different models for the interpretation of the $X(4140)$ and $X(4160)$ resonances. The first model considers that there only exists one wide $X(4140)$ state while the other claims the existence of a narrow $X(4140)$ resonance plus a $X(4160)$ one. Consequently, in this latter case, due to the nature of the $X(4160)$ as a $D^{*}_{s}\bar{D}_{s}^{*}$ [60] bound system, there appears a cusp in the $J/\psi~{}\phi$ mass spectrum near the threshold of this channel $(4224~{}{\rm MeV})$. Therefore, if this cusp is detected, it would be a clear signal that the two-resonance model is the one that represents the nature of the $X$ resonances, and also that the $X(4160)$ is a molecular state.
@@
The $X(4274)$, $X(4350)$ and the $X(4500)$ resonances may also leave a signal at higher $J/\psi$ invariant masses in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay, but they have not been the object of study in the present work.
@@
In the $J/\psi~{}\Xi$ spectrum, we studied the possibility of detecting the double strange pentaquark with hidden charm, which we introduce as a Breit-Wigner using the parameters obtained from the coupled channel approach developed in this work. This permits us to keep the model simple and implement variations on the parameters easily. The results are very promising, since if the mass of this state lies in the region $4580-4680~{}{\rm MeV}$ it has a good chance to be experimentally detected in the spectrum of $J/\psi~{}\Xi$ pairs produced in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay. We have also shown that, in the case of a single wide $X(4140)$ model, the pentaquark signal interferes positively with the background, while, for the two $X$ model, this interference is negative. This potentially allows one to differentiate between these two situations once the corresponding spectrum will be measured.